### v1

# Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization

Abstract
Realistic user simulation is crucial for training and evaluating task-oriented dialogue (TOD) systems, yet creating simulators that accurately replicate human behavior remains challenging. A key property of effective simulators is their ability to expose failure modes of the systems they evaluate. We present an adversarial training framework that iteratively improves user simulator realism through a competitive dynamic between a generator (user simulator) and a discriminator. Applied to mental health support chatbots, our approach demonstrates that fine-tuned simulators dramatically outperform zero-shot base models at surfacing system issues, and adversarial training further enhances diversity, distributional alignment, and predictive validity. The resulting simulator achieves a strong correlation between simulated and real failure occurrence rates across diverse chatbot configurations while maintaining low distributional divergence of failure modes. Discriminator accuracy decreases drastically after three adversarial iterations, suggesting improved realism. These results provide evidence that adversarial training is a promising approach for creating realistic user simulators in mental health support TOD domains, enabling rapid, reliable, and cost-effective system evaluation before deployment.
Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization
Ziyi Zhu1, Olivier Tieleman1, Caitlin A. Stamatis1, Luka Smyth1, Thomas D. Hull1, Daniel R. Cahn1, Matteo Malgaroli2 1Slingshot AI  2Department of Psychiatry, NYU School of Medicine Correspondence: ziyi@slingshotai.com
1 Introduction
Task-oriented dialogue (TOD) systems guide users through multi-turn conversations to achieve specific goals (schatzmann2006survey). The development of large language model (LLM) based TOD systems has created an urgent need for robust evaluation methodologies that go beyond single-turn metrics. Multi-turn task-oriented conversations involve complex dynamics where system quality emerges from sustained interaction, user satisfaction accumulates over time, and task success depends on maintaining context across extended dialogue (ouyang2022training; hong2023zeroshotgoaldirecteddialoguerl). However, traditional evaluation through live user testing is expensive, time-consuming, and exposes users to potentially suboptimal system behavior.
User simulation offers a compelling alternative: synthetic agents that replicate human conversational behavior enable rapid iteration, systematic exploration of edge cases, and controlled experimentation (verma2022chai; yu2023promptbased). At the current state of the art, realistic user simulators have become essential for training language models through reinforcement learning (RL), where they serve as the environment for policy optimization (snell2023offlinerlnaturallanguage; zhou2024archertraininglanguagemodel). Recent advances in multi-turn dialogue optimization have demonstrated the potential of RL methods to improve long-term conversation quality (shani2024multiturnreinforcementlearningpreference; hong2024interactivedialogueagents). However, creating simulators that faithfully reproduce the nuanced, diverse, and goal-directed nature of human conversational behavior presents significant challenges. Simulators that fail to reproduce realistic user behavior cannot reliably expose system failure modes during development, leading to policies that appear effective in simulation but perform poorly with actual users.
In this work, we address the challenge of creating realistic user simulators through adversarial training. Inspired by Generative Adversarial Networks (GANs) (goodfellow2014generative), we train a discriminator to distinguish between simulated and real user sessions, then use the discriminator feedback to refine the simulator through Direct Preference Optimization (DPO) (rafailov2024direct). This creates a competitive dynamic where the simulator learns to fool an increasingly sophisticated discriminator, automatically identifying and correcting the most prominent simulation artifacts.
We evaluate our approach in the context of Ash [^1] https://www.talktoash.com/, a mental health support chatbot designed by Slingshot AI—a complex TOD domain where sustained conversation quality has direct implications for user wellbeing. Our experimental results demonstrate several key findings:
  • Fine-tuning dramatically outperforms zero-shot: Domain-specific fine-tuning is essential for capturing realistic user behavior distributions in TOD, including negative and resistant patterns that zero-shot models fail to exhibit.
  • Adversarial training enhances realism: Three iterations of adversarial refinement progressively recover lexical diversity suppressed by supervised learning, increase distributional similarity, and dramatically reduce discriminator accuracy from near-perfect detection to near-random performance.
  • Strong predictive validity: The best adversarial simulator achieves strong correlation with real model performance and successfully replicates A/B test results on conversation models, enabling reliable offline TOD evaluation before deployment.
These contributions provide evidence that adversarial training is a promising methodology for creating realistic user simulators in the mental health support TOD domain, with potential applications to RL-based policy optimization and reliable offline evaluation. Future work will validate the approach across other TOD domains and integrate simulators into RL training loops.
2 Related Work
2.1 Multi-Turn Dialogue Optimization
While RLHF (ouyang2022training) demonstrated that language models can be aligned with human preferences, it typically optimizes single-turn responses rather than multi-turn dialogue outcomes. DPO (rafailov2024direct) offers an alternative to RLHF by directly optimizing the policy without explicit reward modeling, simplifying the training pipeline while maintaining effectiveness.
Optimizing dialogue systems for long-term conversation quality requires methods that go beyond single-turn reward modeling. Recent approaches include training on imagined conversations and hindsight regeneration (hong2023zeroshotgoaldirecteddialoguerl; hong2024interactivedialogueagents), token-level policy optimization (shani2024multiturnreinforcementlearningpreference), hierarchical decomposition of planning and execution (zhou2024archertraininglanguagemodel), and offline RL methods for language generation (snell2023offlinerlnaturallanguage). These methods critically depend on the quality of the training environment, with realistic user simulators being essential for both policy optimization and evaluation (nam2025efficient).
2.2 User Simulation for Task-Oriented Dialogue
User simulation has been central to TOD research since early agenda-based systems that represented behavior through goal-directed dialogue acts (schatzmann2006survey; schatzmann2007agenda). Neural approaches later enabled more flexible simulation by capturing dialogue history without rigid structures (elasri2016usersim; kreyssig2018neural), demonstrating that policies trained on neural simulators could achieve higher success rates with real users. Recent work on mental health support benchmarking (pombal2024mindeval) demonstrates that LLMs struggle with longer interactions and patients presenting severe symptoms, underscoring the need for realistic simulators that capture the full complexity of human emotions and behaviors.
LLM-Based Simulation: Large language models have transformed user simulation through in-context learning and prompting approaches (li2022dialogic; davidson2023prompting), achieving goal success rates comparable to human-bot interactions with greater linguistic diversity and minimal seed dialogues. However, prompting-only approaches suffer from hallucinations and goal drift. Fine-tuning addresses these limitations through various techniques, including reducing behavioral inconsistencies, joint policy and NLG optimization for domain transfer, and dual-model architectures with chain-of-thought reasoning (sekulic2024daus; lin2022gentus; luo2024duetsim).
RL-Based Simulator Training: While supervised learning dominates simulator training, reinforcement learning approaches optimize for task-level objectives rather than imitation alone. Multi-agent RL enables joint system-user training (takanobu2020guided; liu2018dialogue), while adversarial training creates competitive dynamics between simulators and discriminators (keizer2023adversarial; li2017adversarial). The adversarial approach has demonstrated substantial improvements, with discriminator-based rewards being "more general, natural and effective" than hand-crafted metrics. Our work extends this adversarial paradigm to LLM-based simulators, using DPO rather than policy gradient methods to leverage discriminator feedback while avoiding reward hacking.
Evaluation Methodology: Traditional TOD evaluation relies on task-level metrics such as success rate and slot error rate measured on standard benchmarks like MultiWOZ (budzianowski2018multiwoz), SGD (rastogi2020schema), and tau-bench (yao2024taubench), which effectively capture transactional outcomes but provide limited insight into conversational quality. The poor correlation between surface-level metrics like BLEU/ROUGE and human judgments (liu2016not) motivates more sophisticated evaluation approaches that consider behavioral fidelity and distributional alignment (nam2025efficient). Our work advances this evaluation paradigm through three key contributions: (1) explicit modeling of conversational feature occurrence rate distributions using KL divergence to measure alignment on fine-grained failure modes rather than binary task success, (2) validation of predictive utility through model A/B test replication beyond correlation analysis, and (3) discriminator accuracy as a direct measure of distributional realism.
3 Methodology
3.1 Problem Formulation
We formalize user simulation as a conditional generation task. Given conversation history $h_{t}=\{(u_{1},a_{1}),...,(u_{t-1},a_{t-1}),a_{t}\}$ consisting of previous user utterances $u_{i}$ and agent responses $a_{i}$, plus user context $c$ (e.g., interaction history and background for the user), the simulator generates the next user utterance $u_{t}\sim\pi_{\theta}(u_{t}|h_{t},c)$.
We train $\pi_{\theta}$ to generate simulated conversations indistinguishable from real user conversations. To this end, we bootstrap an initial simulator from real conversations, construct prompts with appropriate context, and iteratively alternate between training discriminators to detect simulated conversations and refining simulators to evade detection.
3.2 Context Construction for User Simulation
A critical design choice for realistic user simulation is the construction of context $c$ that conditions the simulator’s behavior. Effective context should provide: (1) role and behavioral guidelines through system instructions, (2) user background and goals, (3) conversational history to maintain consistency, and (4) session-specific grounding to anchor the simulator to target conversation characteristics. The specific context structure depends on the domain—transactional TOD systems might include task schemas and slot values, while open-ended domains might emphasize user profiles and session summaries. Our mental health support chatbot implementation (detailed in Appendix A) uses a hierarchical structure spanning user profiles, multi-session summaries, and current session grounding. As we demonstrate in Section 5, rich context grounding enables fine-tuned simulators to cluster tightly with their corresponding real sessions in embedding space, while zero-shot models without such grounding form distinct clusters, indicating systematic behavioral drift.
3.3 Base Simulator Training
We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations. Starting from Llama-3.3-70B-Instruct as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(h,c,u)\sim\mathcal{D}}[\log\pi_{\theta}(u|h,c)]$ (1)
3.4 Discriminator Training
To measure simulator realism, we train a discriminator $D_{\phi}$ to distinguish between simulated and real user sessions. We implement the discriminator as a token classification model built on the Llama-3.1-8B-Instruct backbone. The discriminator processes the full conversation and predicts after each user message whether the session is real or simulated:
$p_{t}=D_{\phi}(\text{real}|h_{t},c)$ (2)
We use causal attention in the discriminator, meaning predictions are made based on the current message and all previous messages in the session. This design choice is essential because any individual message could plausibly be either real or simulated when taken out of context. The discriminator’s effectiveness stems from detecting distributional patterns across the conversation: certain phrases repeated multiple times, unnatural topic progressions, or behavioral patterns that diverge from real user dynamics. Without conversational context, the discrimination task becomes noisy and ill-defined, as isolated messages lack the sequential dependencies that reveal simulation artifacts.
We train $D_{\phi}$ on a balanced dataset of real and simulated sessions using binary cross-entropy loss:
$\mathcal{L}_{\text{disc}}=-\mathbb{E}_{(h,c,y)}[y\log p+(1-y)\log(1-p)]$ (3)
where $y=1$ for real sessions and $y=0$ for simulated sessions. A high discriminator accuracy indicates the simulator has not achieved realistic behavior; our goal is to train simulators that reduce discriminator accuracy toward random guessing.
In our adversarial training approach, each new discriminator is trained exclusively on data from the most recent simulator version. This ensures the preference pairs we generate for DPO reflect the current policy’s behavior rather than artifacts from earlier iterations. Research on preference optimization supports this on-policy approach: on-policy preference pairs exhibit stronger contrastiveness and clearer preference signals (zhao2024reflective), online methods require only “partial coverage” compared to the strict “global coverage” needed by offline methods (song2024importance), and on-policy sampling efficiently concentrates probability mass on high-reward responses (tajwar2024preference).
3.5 Adversarial Training Framework
Our adversarial training process follows an iterative refinement loop, illustrated in Figure 1. Algorithm 1 provides the detailed procedure.
Algorithm 1 Adversarial User Simulator Training
1: Initialize user simulator $\pi_{\theta}$ from SFT
2: Initialize discriminator $D_{\phi}$
3:for iteration $k=1,2,$... do
4:  Generate simulated sessions $\mathcal{S}_{k}$ using $\pi_{\theta}$
5:  Train discriminator $D_{\phi}$ on real vs. simulated sessions from $\pi_{\theta}$
6:  Compute rewards $r_{t}$ and identify high/low reward messages in $\mathcal{S}_{k}$
7:  Sample alternative responses at critical points to create preference dataset $\mathcal{D}_{\text{pref}}$
8:  Retain only those pairs satisfying $r_{\text{chosen}}>0$ and $r_{\text{rejected}}<0$
9:  Update simulator using DPO: $\pi_{\theta}\leftarrow\text{DPO}(\pi_{\theta},\mathcal{D}_{\text{pref}})$
10:end for
Figure 1: Overview of the adversarial training process. The simulator generates sessions that are evaluated by the discriminator to distinguish real from simulated conversations. Discriminator feedback guides preference dataset generation, which is then used to refine the simulator through DPO, creating an iterative competitive dynamic that improves realism.
Optimization Method: We use DPO rather than RLHF methods (schulman2017proximal; shao2024deepseekmath) because discriminator rewards are inherently policy-dependent and vulnerable to reward hacking. With RLHF, the simulator would continuously optimize against a fixed reward model, eventually exploiting patterns that score highly but don’t reflect genuine realism. Preventing this would require retraining the discriminator throughout optimization—a substantial computational overhead. DPO avoids this problem: we generate preference pairs once per iteration using the current discriminator, then optimize against this fixed dataset. Preferences grounded in comparisons between specific alternatives are more robust than absolute reward values, and the discrete iteration structure naturally prevents reward hacking.
Reward Signal: For each user message in a simulated session, we compute a reward based on how the message changes the discriminator’s prediction:
$r_{t}=\log\left(\frac{p_{t}}{1-p_{t}}\right)-\log\left(\frac{p_{t-1}}{1-p_{t-1}}\right)$ (4)
We use log-odds rather than raw probabilities to reflect the exponentially increasing difficulty of fooling the discriminator at higher confidence levels. Higher rewards indicate messages that make the session appear more realistic.
Preference Dataset Generation: At messages with the highest and lowest rewards (critical moments where the discriminator most clearly detects unrealism), we sample 8 alternative responses. These alternatives are ranked by their discriminator-based rewards to create preference pairs $(u_{\text{chosen}},u_{\text{rejected}})$ where the chosen response has the highest reward.
Dataset Filtering: After generating preference pairs, we filter to retain only samples where $r_{\text{chosen}}>0$ and $r_{\text{rejected}}<0$. This selects high-confidence pairs with maximum learning signal while preventing over-correction.
DPO Update: We train the new simulator using Direct Preference Optimization (rafailov2024direct), which directly optimizes the policy to prefer higher-reward responses. We use the previous iteration’s simulator as the reference policy $\pi_{\text{ref}}$. Defining the implicit reward $r_{\theta}(h,c,u)=\log\frac{\pi_{\theta}(u|h,c)}{\pi_{\text{ref}}(u|h,c)}$, the DPO loss is:
$\begin{split}\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(h,c,u_{c},u_{r})\sim\mathcal{D}_{\text{pref}}}\Big[\log\sigma\Big(\beta\big(r_{\theta}(h,c,u_{c})\\ \quad-r_{\theta}(h,c,u_{r})\big)\Big)\Big]\end{split}$ (5)
where $u_{c}$ and $u_{r}$ are chosen and rejected responses, $\beta$ is the temperature parameter, and $\sigma$ is the sigmoid function.
4 Experimental Setup
4.1 Domain and Data
We evaluate our methodology in the context of mental health support—a complex task-oriented dialogue domain where users pursue mental health goals (emotional regulation, insight development, behavioral change) through extended multi-turn conversations. Unlike traditional TOD domains with discrete transactional goals, mental health support offers distinct advantages for studying realistic user simulation: rich emotional dynamics and behavioral complexity, multi-session continuity, diverse naturalistic failure modes beyond binary task success, and production-scale data capturing rare but critical patterns like crisis expressions and resistance to engagement.
Our dataset consists of anonymized real user conversations with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).
4.2 Model Configurations
We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (hu2022lora) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects. We perform adversarial training on the final checkpoint (160K), producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations. For comparison, we evaluate three zero-shot base models: GPT-4o, Kimi-K2-Instruct, and Llama-3.3-70B-Instruct. Training hyperparameters are provided in Appendix C.
The discriminators (UserDisc-base, UserDisc-it1, UserDisc-it2, UserDisc-it3) are token classification models trained from the Llama-3.1-8B-Instruct backbone.
For each configuration, we generate 2,000 simulated sessions: 1,000 unique session contexts drawn from real user profiles and conversation histories × 2 simulations per context, with early stopping when conversations naturally conclude. Our context construction for user simulation follows the hierarchical structure described in Appendix A, including user profiles, multi-session history, and current session grounding. To isolate simulator quality effects, we use a fixed set of 14 chatbot models to simulate conversations.
4.3 Evaluation Metrics
4.3.1 Linguistic Features
For all linguistic features, we calculate metrics for each session individually and report distributional statistics (mean, median, standard deviation) across sessions to capture both typical behavior and variability.
Word Entropy: Vocabulary diversity using Shannon entropy: $H=-\sum_{w}p(w)\log p(w)$, where $p(w)$ is the probability of word $w$ in the session’s vocabulary. Higher entropy indicates more diverse word usage.
Type-Token Ratio (TTR): The ratio of unique words (types) to total words (tokens) in a session, measuring lexical diversity. TTR ranges from 0 to 1, with higher values indicating greater vocabulary variety.
Bigram Diversity: The ratio of unique word pairs (bigrams) to total bigrams in a session, measuring phrase-level variety and reducing sensitivity to repetitive phrasing.
4.3.2 Behavioral Metrics
User Action Distribution: Conversations are annotated with mental health-relevant user actions using GPT-5-nano to classify each individual user message according to our action taxonomy (see Appendix D for example categories). Since individual messages can have multiple actions, the action distributions are not normalized (probabilities do not sum to 1).
Issue Rate Analysis: A critical requirement for realistic user simulation is the ability to expose chatbot failure modes. We evaluate simulators by measuring how well they replicate the rate and distribution of problematic conversational features (“issues”) that chatbots exhibit with real users. Issues are identified using GPT-5 with low reasoning effort to analyze entire sessions. The analyzer identifies issues from a detailed taxonomy (35+ categories including invalidating responses, unclear wording, hallucinations, etc.) and highlights the specific messages where each issue occurred, enabling proper credit assignment (see Appendix E for example categories). We compare issue rates and category distributions between simulated and real conversations to validate that our simulator accurately reflects the chatbot’s real-world performance and can reliably identify system weaknesses before deployment. Since we examine proportions of issue categories (which form a valid probability distribution that sums to 1), we use KL divergence to measure distributional alignment:
$\text{KL}(p\|q)=\sum_{i}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)$ (6)
where $p$ is the real distribution and $q$ is the simulated distribution.
Similar to the discriminator, the issue analyzer examines whole sessions rather than isolated messages. This design is essential because messages are only problematic when considered in context with the user’s situation and their subsequent replies. The session-level approach makes realistic user simulation critical for accurate evaluation—the simulator must produce contextually appropriate responses that reveal when the chatbot’s messages fail to address the user’s actual needs. This enables numerous evaluation dimensions that would be impossible with isolated message analysis: detecting circular conversations, identifying premature topic changes, recognizing when the chatbot ignores user requests, and assessing whether mental health progress occurs naturally over the conversation arc.
4.3.3 Discriminator Performance
We measure discriminator accuracy using fine-tuned discriminators trained on balanced datasets of real and simulated sessions from their corresponding simulator versions. Each discriminator is trained under identical conditions—1,600 training sessions and 400 validation sessions, with the same learning rate and number of epochs—ensuring that accuracy differences reflect simulator quality rather than training differences. Lower discriminator accuracy indicates a more realistic simulator, with the target being 50% (random guessing).
5 Results
5.1 Fine-tuned vs. Zero-shot Comparison
Fine-tuned simulators dramatically outperform zero-shot models across all metrics. t-SNE visualization (Appendix B.1) shows that embeddings of conversations simulated by base models form distinct clusters separate from their real counterparts. The fine-tuned simulators’ outputs, on the other hand, form tight clusters defined by shared context rather than by the generation process, demonstrating effective grounding. Zero-shot models produce unrealistically compliant users, as evidenced by significantly lower issue rates (Figure 2) and severe overrepresentation of positive mental health indicators (Appendix B.3).
Table 1: Linguistic diversity features comparing real users with simulators. Entropy measures vocabulary diversity; TTR (type-token ratio) measures lexical diversity; Diversity measures bigram diversity. Values closer to real users indicate better linguistic realism. Best fine-tuned values (closest to real users) are in bold.
Simulator Entropy TTR Bigram
Real Users 6.60 0.483 0.861
GPT-4o 7.12 0.408 0.837
Kimi-K2-Instruct 7.65 0.416 0.851
Llama-3.3-70B-Instruct 7.02 0.268 0.651
UserSim-10K-base 6.14 0.406 0.756
UserSim-20K-base 6.23 0.422 0.782
UserSim-40K-base 6.14 0.430 0.782
UserSim-80K-base 6.15 0.409 0.759
UserSim-160K-base 6.18 0.382 0.728
UserSim-160K-it1 6.38 0.471 0.839
UserSim-160K-it2 6.52 0.439 0.821
UserSim-160K-it3 6.51 0.461 0.841
As shown in Table 1, which compares linguistic features across simulators, some base models achieve diversity metrics closer to real users than fine-tuned models (e.g., Kimi-K2-Instruct’s bigram diversity of 0.851 vs. real users’ 0.861) as supervised fine-tuning exhibits non-monotonic effects on lexical diversity. Bigram and TTR initially improve from 10K-40K samples as the model learns to use conversational context to improve response diversity. However, mode collapse emerges from 40K-160K samples as continued training learns common patterns at the expense of rare variations. For behavioral alignment, fine-tuned simulators dramatically outperform zero-shot base models, achieving total BCE losses of .159-.160 compared to .187-.207 for zero-shot models (see Appendix B.3). To evaluate how well simulators surface chatbot issues, we measure issue rates (percentage of chatbot messages containing issues) using GPT-5 to analyze complete sessions. Figure 2 shows that fine-tuned simulators produce rates of 3.07-3.31%, closely matching real conversations (3.02%), while zero-shot models show substantially lower rates of 0.98-2.54%. UserSim-160K-it2 achieves the closest calibration to the real rate (3.07%). Detailed issue category discrepancies are provided in Appendix B.2.
Figure 2: Issue rate comparison across simulated and real sessions. Error bars represent standard error. Zero-shot base models significantly underestimate issue rates, while fine-tuned models achieve realistic calibration.
5.2 Adversarial Training Improvements
Three iterations of adversarial training produced progressive improvements:
Table 2: Discriminator performance (accuracy, F1, MCC) at different conversation turns. MCC (Matthews Correlation Coefficient) is a measure of the correlation between the observed and predicted classifications. Lower performance indicates more realistic simulation. Best results (closest to target) are in bold.
Turn 1 Turn 5 Turn 10 Turn 25 Overall
Discriminator Acc F1 MCC Acc F1 MCC Acc F1 MCC Acc F1 MCC Acc F1 MCC
UserDisc-base 62.0 80.1 24.3 82.3 81.6 64.7 90.2 88.6 80.1 96.1 93.9 91.1 87.9 85.9 75.4
UserDisc-it1 52.8 53.8 5.5 70.3 69.3 40.6 79.5 76.0 58.3 87.4 77.4 68.7 76.5 71.4 51.5
UserDisc-it2 53.8 51.3 8.2 67.3 65.3 34.7 76.2 71.6 51.3 90.1 86.5 79.2 74.0 69.1 47.2
UserDisc-it3 52.5 60.6 5.5 64.0 64.4 28.0 68.1 59.5 33.9 84.5 78.9 67.3 69.5 63.9 37.6
Linguistic Diversity: Adversarial training successfully addresses mode collapse by recovering lexical diversity while maintaining behavioral alignment. UserSim-160K-it3 progressively approaches real user levels across all diversity metrics, demonstrating that adversarial training addresses a fundamental limitation of supervised fine-tuning. However, oscillatory behavior across iterations indicates the need for careful hyperparameter tuning. We employ conservative dataset filtering (retaining only high-confidence preference pairs) to mitigate over-correction and maintain distribution stability.
Distributional Alignment: Adversarial training produces modest improvements in action distribution alignment (see Appendix B.3), with progressive refinement across iterations. The limited magnitude of improvements suggests the base fine-tuned model already achieved strong behavioral alignment.
Discriminator Fooling: Adversarial training dramatically reduced discriminator accuracy across conversation turns (Table 2). All discriminators were trained under identical conditions with on-policy optimization—each trained exclusively on sessions from its corresponding simulator version. The dramatic performance degradation confirms that adversarial iterations remove systematic artifacts and produce strong distributional alignment throughout full sessions.
Predictive Validity: To evaluate how well simulators enable reliable offline evaluation, we measure correlation between simulated and real issue rates across diverse chatbot models, and KL divergence between simulated and real issue category distributions. Table 3 shows that adversarial training produces substantial improvements in both metrics. UserSim-160K-it2 achieves the strongest correlation ($r=0.818$), while UserSim-160K-it3 achieves the lowest distributional divergence (KL = 0.148). This provides compelling evidence that adversarially trained simulators are significantly more realistic and effective at identifying chatbot problems in offline testing compared to their base fine-tuned counterparts. Notably, aggregating predictions across all simulator models achieves a strong correlation ($r=0.801$) and low distributional divergence (KL = 0.152), suggesting that ensembling multiple simulators could provide even more robust evaluation.
Table 3: Correlation with real issue rates and KL divergence of issue category distributions. Correlation computed across 14 diverse chatbot configurations. Aggregate combines UserSim-160K-base, it1, it2, and it3. Best values are in bold.
Simulator $r$ $p$ KL Div
GPT-4o 0.484 0.079 1.566
Kimi-K2-Instruct 0.564 0.036 0.453
Llama-3.3-70B-Instruct 0.629 0.016 0.542
UserSim-160K-base 0.528 0.052 0.671
UserSim-160K-it1 0.685 0.007 0.630
UserSim-160K-it2 0.818 0.0003 0.318
UserSim-160K-it3 0.781 0.001 0.148
Aggregate 0.801 0.0006 0.152
A/B Test Replication: Beyond aggregate correlation, we validate predictive utility by replicating a recent model A/B test. All three models exhibited consistent directional trends between simulation and real data, and no model that failed in simulation improved in the wild (Appendix B.5). Combined with strong correlation and distributional alignment, these results validate the simulator’s utility for reliable offline evaluation.
6 Discussion
Our results establish several important findings for user simulator development:
Firstly, zero-shot models exhibit severe distributional biases, producing unrealistically compliant users who systematically underrepresent negative and resistant behaviors. This confirms that generic language models, despite their capabilities, cannot reliably replicate domain-specific user behavior without exposure to real conversational patterns.
While supervised fine-tuning achieves strong behavioral alignment, it suppresses lexical diversity through mode collapse. Adversarial refinement recovers this diversity while maintaining behavioral fidelity—successfully producing simulators that fool increasingly sophisticated discriminators without sacrificing predictive utility. This demonstrates that realism requires both distributional accuracy and surface-level variation.
Lastly, unlike traditional metrics that measure task success or correlation, discriminator accuracy directly quantifies distributional realism by testing whether conversations are distinguishable from real interactions. The dramatic accuracy degradation across adversarial iterations provides evidence of improved simulation fidelity that complements task-level validation.
7 Conclusion
We have presented an adversarial training framework for creating realistic user simulators that can reliably expose failure modes in multi-turn task-oriented dialogue systems. Applied to mental health support—a complex TOD domain where users pursue mental health goals through extended goal-directed conversations—our approach demonstrates that domain-specific fine-tuning combined with adversarial refinement produces simulators that closely replicate real user behavior across linguistic, semantic, and behavioral dimensions.
Our key contributions are: (1) demonstration that fine-tuned models vastly outperform zero-shot base models for behavioral fidelity, (2) an adversarial training method that recovers lexical diversity suppressed by supervised learning while improving distributional alignment and predictive validity, and (3) comprehensive validation that adversarially-trained simulators reliably predict real model performance, achieving strong correlation with real-world results and successfully replicating A/B test trends.
These results provide evidence that adversarial training is a promising methodology for creating realistic user simulators in mental health support TOD domains. By enabling rapid iteration and systematic exploration of system failure modes before deployment, such simulators can reduce development time, protect users from suboptimal experiences, and potentially provide safe environments for reinforcement learning based optimization of conversation quality.
Limitations
While our approach demonstrates significant improvements in simulator realism, several limitations remain that warrant future investigation:
Mode Collapse: While adversarial training substantially improves lexical diversity, fine-tuned simulators still show lower phrase-level variety than real users. Techniques like controlled generation or mixture-of-experts architectures might further improve diversity preservation.
Domain Transfer: Our evaluation focuses on mental health support—a complex, open-ended TOD domain. Validating the methodology in traditional TOD domains (restaurant booking, customer service) and other complex domains (tutoring, technical support) would establish broader applicability and identify domain-specific challenges.
RL Training Integration: While our simulators enable valid offline evaluation, integrating them directly into RL training loops for TOD policy optimization remains future work. Recent work on RL-based tutoring (nam2025efficient) demonstrates that simulator quality is critical for reliable RL-based dialogue optimization, underscoring the importance of adversarial training for creating simulators that support both evaluation and training.
Ethical Considerations
All data used in this research were collected from real user conversations with our mental health chatbot with explicit user consent. All personally identifiable information was removed through anonymization procedures in accordance with data protection policies and ethical guidelines for human subjects research. However, due to the sensitive nature of mental health data, we are unable to share the datasets used in this research, despite our desire to promote reproducibility and open science.
Our simulators are designed for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences. We note that simulators trained on real conversations may perpetuate biases present in the original dataset; applications should include appropriate oversight and ethical review. Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions or edge cases.
Acknowledgments
We thank the reviewers for their constructive feedback. We are grateful to the Slingshot AI team for their valuable discussions, ideas, and foundational work that made this research possible. Special thanks to Jinghong Chen for his help in reviewing and improving this work. Finally, we extend our deepest gratitude to the users of Ash, our mental health support chatbot, who opted in to provide the data used for this study.
Appendix A Context Construction for Mental Health Support Chatbot
This appendix describes the specific context structure used for our mental health support chatbot user simulator. While the general principles (role definition, user background, conversational history, session grounding) apply broadly to TOD systems, the specific implementation reflects the requirements of mental health support.
Our context construction follows a hierarchical structure with the following components:
System Instruction: Explicit instructions defining the simulator’s role as a user in mental health support conversations, including guidelines for consistency and realistic behavior.
User Profile: Demographics and onboarding responses establishing presenting concerns, goals, and background information. This captures the user’s initial state and motivations for seeking support.
Long-term Context: Aggregated information about the user’s mental health support journey, including recurring themes, progress over time, and persistent concerns that span multiple sessions. This enables the simulator to maintain consistency across extended timeframes.
Previous Session Summaries: Summaries of 3-5 prior sessions capturing key topics discussed, insights gained, and mental health support interventions attempted. This provides cross-session continuity without requiring the full message history.
Current Session Summary: A summary of the current session being simulated, extracted from the original conversation. This grounds the simulator in the specific topics and emotional tone of the target session, reducing aimless or disengaged responses. This session-specific grounding is critical for the simulator to cluster tightly with corresponding real sessions.
Raw Message History: Message-by-message conversation history from recent sessions, providing detailed behavioral context and immediate conversational flow.
The complete context is truncated to 15,000 tokens, prioritizing recent session content and current session summary over distant history. All context components are automatically generated using production usage—user profiles from onboarding and session summaries after each conversation—with approximately 100,000 user profiles available for training and evaluation. This enables our approach to scale naturally with system usage without requiring manual annotation or synthetic data generation.
Appendix B Detailed Analysis
This appendix provides detailed visualizations and analysis supporting the main results, including embedding visualizations, correlation analysis, and action distribution comparisons.
B.1 Embedding Visualizations
We embed all sessions using OpenAI’s text-embedding-3-large, including up to 50 messages per session within the model’s token limit. We compute t-SNE projections with perplexity 30 on all simulated and real sessions, then visualize 50 randomly sampled sessions for clarity.
Figure 3: t-SNE visualization of session embeddings based on all messages. Production sessions are marked with stars. Base models (circles) form distinct clusters separate from real conversations, while fine-tuned simulators (squares) cluster tightly around their corresponding real sessions, demonstrating effective context grounding.
Figure 3 shows the t-SNE visualization of session embeddings. Each base model (GPT-4o, Kimi-K2-Instruct, Llama-3.3-70B-Instruct) forms a distinct cluster (circles) separate from real sessions (stars), indicating systematic behavioral differences. In contrast, fine-tuned simulators (squares) cluster tightly around their corresponding real sessions, demonstrating that rich context grounding enables simulators to maintain topical coherence and behavioral fidelity.
B.2 Issue Category Distribution
Figure 4: Issue category discrepancy between UserSim-160K-it3 and real conversations. Left: categories over-represented in simulation (positive delta indicates simulation detects more). Right: categories under-represented in simulation (negative delta indicates real sessions have more). Deltas shown as percentage point differences.
Figure 4 shows the discrepancy between simulated and real session issue category distributions. The simulator over-represents certain categories, including “Chatbot contradicts itself” ($\sim$1.5 percentage points), “Chatbot forgets in context” ($\sim$1.25 percentage points), and “unclear or confusing wording”—likely because simulated users engage more persistently with problematic chatbot behaviors, triggering repeated issues. Conversely, the simulator under-represents categories like “duplicate chatbot messages” ($\sim$-2.5 percentage points), “pushing techniques after user request to stop” ($\sim$-2.5 percentage points), and “invalidating or dismissive framing,” which require specific conversational histories that may be less common in simulation. The relatively small magnitude of most discrepancies (within $\pm$2.5 percentage points) indicates reasonable distributional alignment overall.
B.3 User Action Distributions
Figure 5: User action distributions comparing simulators to real data. Left: negative mental health indicators (Bad). Right: positive mental health indicators (Good). Error bars represent standard error. Zero-shot base models (GPT-4o, Kimi-K2-Instruct, Llama-3.3-70B-Instruct) show severe distributional misalignment, while fine-tuned simulators closely match real distributions.
Figure 5 compares user action distributions across simulators. For negative indicators (left), zero-shot models dramatically underestimate rates of challenging behaviors: GPT-4o shows near-zero “Aggressive to Chatbot” while real users exhibit $\sim$0.5%, and all base models underrepresent “Asks for Clarity” and “Disagrees with Chatbot.” This confirms that zero-shot models produce unrealistically compliant users.
For positive indicators (right), zero-shot models exhibit the opposite bias—substantially overrepresenting “Change Talk” ($\sim$40% vs. $\sim$15% in real data) and “Agrees with Chatbot.” This creates an artificially positive mental health support environment that fails to challenge the chatbot or surface issues. Fine-tuned simulators (UserSim variants) achieve close alignment with real sessions across both indicator types, with adversarial training iterations showing incremental improvements in matching real distributions.
We use Binary Cross-Entropy (BCE) to measure distributional alignment:
$\text{BCE}(p,q)=-\sum_{i}[p_{i}\log(q_{i})+(1-p_{i})\log(1-q_{i})]$ (7)
where $p$ is the real distribution and $q$ is the simulated distribution. BCE is appropriate for such multi-label distributions where action probabilities do not sum to 1, while KL divergence requires proper probability distributions.
Table 4: BCE loss from real user action distributions by category. Lower is better. Fine-tuned simulators (.159-.160) substantially outperform zero-shot base models (.187-.207), demonstrating better behavioral alignment. Best values are in bold.
Simulator Total Bad Ind. Good Ind. Moderators User Feat.
GPT-4o .18662 .08772 .18573 .10637 .29522
Kimi-K2-Instruct .19116 .06451 .20626 .10414 .30948
Llama-3.3-70B-Instruct .20721 .06269 .24554 .10466 .32205
UserSim-160K-base .15866 .05994 .13680 .10467 .28056
UserSim-160K-it1 .15962 .05975 .13731 .10434 .28333
UserSim-160K-it2 .15861 .05968 .13697 .10401 .28065
UserSim-160K-it3 .15866 .05970 .13658 .10409 .28114
B.4 Correlation with Production
Figure 6: Correlation between simulated and real issue rates for UserSim-160K-it2 across 14 chatbot model configurations spanning multiple model families (Qwen-based, Llama-based, GPT-based). Error bars represent standard error. Strong correlation ($r=0.818$, $p=0.0003$) validates the simulator’s ability to predict relative model performance before production deployment.
Figure 6 shows the correlation between simulated and real issue rates across 14 diverse chatbot configurations. The Pearson correlation coefficient $r=0.818$ indicates a strong linear relationship, with the simulator correctly ranking model quality. The p-value ($p=0.0003$) represents the probability of observing this correlation by chance under the null hypothesis of no correlation. It is calculated using a t-test: $t=r\sqrt{(n-2)/(1-r^{2})}$ with n2n-2 degrees of freedom.
The adversarial training process substantially improves correlation with real issue rates, as shown in Table 3. UserSim-160K-it2 achieves $r=0.818$ compared to UserSim-160K-base’s $r=0.528$, demonstrating the value of adversarial refinement. However, we note that several models achieve strong correlations even without adversarial training. This suggests that for rolled-out evaluation or RL training applications, ensembling multiple user simulator models could reduce noise and improve stability. Indeed, aggregating across UserSim-160K-base, it1, it2, and it3 achieves robust correlation ($r=0.801$) with low distributional divergence (KL = 0.152). The plot shown here highlights the best-performing model (UserSim-160K-it2) to demonstrate the upper bound of achievable predictive validity.
B.5 A/B Test Replication
To validate predictive utility beyond aggregate correlation, we replicate a recent A/B test comparing control chatbot models against treatment versions. The treatment models were fine-tuned using DPO on synthetic issue data generated by our simulator, with the goal of reducing issue rates. We selected this experiment because it directly targets issue rates—a significant metric we aim to predict with the simulator—and produced statistically significant results that provide a clear benchmark for replication.
The intervention ultimately failed to achieve its intended goal, with the fine-tuned models actually increasing issue rates rather than reducing them. This negative result makes it an ideal test case: a valid simulator should correctly identify that the fine-tuning applied here is ineffective.
Table 5: Issue rate comparison between simulated and real evaluations with two-tailed significance tests ($\alpha=0.05$). The treatment (fine-tuned with DPO on synthetic issue data to reduce issues) failed and actually increased issue rates. Both simulation and real data show consistent directional trends, validating the simulator’s predictive utility.
Simulation (UserSim-160K-base) Production
Model Control Treatment $z$ $p$ Control Treatment $z$ $p$
Model A (Qwen-based) 2.65% 3.25% 2.83 0.005* 2.71% 3.56% 3.27 0.001*
Model B (Llama-based) 2.42% 2.87% 2.23 0.026* 2.32% 2.64% 1.35 0.178
Model C (GPT-based) 1.78% 2.13% 2.02 0.044* 1.54% 1.42% -0.66 0.511
*Statistically significant at $\alpha=0.05$
Table 5 shows that all three models exhibited consistent directional trends between simulation and real data. Critically, no model that failed in simulation improved in real sessions, demonstrating that the simulator successfully identifies ineffective interventions before deployment.
Appendix C Training Hyperparameters
Table 6 shows the hyperparameters used for supervised fine-tuning (SFT) of the base user simulator and Direct Preference Optimization (DPO) during adversarial training iterations.
Table 6: Hyperparameters for supervised fine-tuning (SFT) and Direct Preference Optimization (DPO).
Hyperparameter SFT DPO
Epochs 1 1
Batch size 1 1
DPO beta ($\beta$) 1
LoRA rank 8 8
LoRA alpha 16 16
LoRA trainable modules all-linear all-linear
Learning rate $5\times 10^{-5}$ $1\times 10^{-5}$
Learning rate scheduler constant constant
Warmup ratio 0 0.1
Max gradient norm 1 1
Weight decay 0 0
Appendix D User Action Categories
This appendix provides example user action categories from our taxonomy used to annotate and analyze user behavior in mental health support conversations. Each user message is classified using GPT-5-nano. Actions are classified as positive mental health indicators (Good), negative indicators (Bad), or user features that should change over time. The full taxonomy contains 20+ categories; we highlight key examples below.
D.1 Example Positive Indicators
  • Gained an insight: User expresses that they learned something new about themselves or about their situation.
  • Change talk: User uses change talk to set a task goal or commit to a healthy action—something specific.
  • Agrees with chatbot: User agrees with an observation, suggestion, or advice that the chatbot gives.
  • Chatbot created change: User identifies something positive that has changed as a result of talking to the chatbot.
D.2 Example Negative Indicators
  • Disagrees with chatbot: User disagrees with an observation, suggestion, or advice that the chatbot gives.
  • Asks for clarity: User asks for the chatbot to clarify what it’s doing or saying.
  • Aggressive to chatbot: User gets angry with the chatbot or is combative or abusive in some way.
  • Chatbot loses user: User is not sure how to respond to the chatbot and responds in a way that suggests they’re disengaging.
D.3 Example User Features
  • General distress: User reports discomfort or suffering that is vague and lacks a specific object or expresses not knowing who they are or what life means to them.
  • Negative affect: User is able to acknowledge a negative emotion and make it explicit—a specific emotion must be identified; general expressions of negative affect fall under general distress.
  • Positive affect: User expresses a positive emotion and makes it explicit—a specific emotion must be identified.
Appendix E Issue Categories
This appendix provides example issue categories from our taxonomy used to evaluate chatbot quality. Issues are identified by GPT-5 analyzing entire conversation sessions with low reasoning effort. For each detected issue, the analyzer highlights the specific affected messages to enable proper credit assignment. The full taxonomy contains 35+ categories organized into six main categories (User Distress, User Confusion, User Frustration, User Losing Trust, Chatbot Violates Policies, App Bugs); we highlight key examples below.
E.1 User Distress Examples
  • Invalidating or dismissive framing: The chatbot uses language that minimizes the user’s struggle or suggests they could simply choose to feel better (e.g., framing complex emotional states as mere choices).
  • Unempathetic responses: The chatbot responds in ways that lack empathy in distressing or painful situations, including being overly brief, blunt, cold, detached, or otherwise failing to appropriately acknowledge or validate the user’s emotional state.
  • Premature wrap-up: The chatbot initiates ending the session or wraps up the conversation when the user has explicitly expressed wanting to continue; it’s clear the user doesn’t want to end; the user is mid-sentence; or at otherwise inappropriate moments, such as immediately after the user has shared something sensitive or vulnerable.
E.2 User Confusion Examples
  • Chatbot hallucinates: The chatbot invents or misremembers information about the user—such as their experiences, feelings, or history—and the user explicitly corrects or objects to the inaccuracy.
  • Unclear or confusing wording: The chatbot says something that is unusually complex, convoluted, or strange, making it difficult for the user to understand the intended meaning or follow the flow of conversation.
  • Chatbot contradicts itself: The chatbot makes statements or suggestions that directly contradict each other in meaning, tone, or advice, creating clear confusion or inconsistency that cannot reasonably be explained as exploring multiple perspectives.
E.3 User Frustration Examples
  • Circular conversations: The chatbot leads the conversation in circles, revisiting the same topics or lines of discussion that have already been addressed, without adding new insight or progress. The issue applies only when the user shows clear signs of frustration or impatience with this repetition.
  • Ignoring user’s request to end session: When the user clearly indicates they need to leave or pause or end the conversation (explicitly or implicitly), the chatbot ignores this and continues asking questions or advancing the discussion instead of acknowledging and closing promptly.
  • Repetitive questions: The chatbot repeats a question or statement that is nearly identical to something it already said earlier in the same session, without any clear reason or therapeutic justification for the repetition.
E.4 User Losing Trust Examples
  • Chatbot gives bad advice: The chatbot gives advice or suggestions that are objectively poor, unhelpful, or clearly irrelevant to the user’s situation—not merely advice that the user disagrees with or finds unappealing.
  • Chatbot is sycophantic: The chatbot responds with excessive, exaggerated, or insincere praise or agreement that goes beyond normal empathy, validation, or encouragement. The issue applies only when the flattery or agreement is so strong or uncritical that it feels unnatural, manipulative, or disingenuous.
E.5 Policy Violation Examples
  • Failure to follow crisis protocol: The chatbot does not provide required crisis resources, encourage immediate human help, or include the mandated disclaimer about AI limitations when suicide or imminent harm is mentioned.
  • Providing medical diagnoses: The chatbot gives specific medical or mental health diagnoses, which it is not qualified to provide as an AI system.
  • Inappropriate self-disclosure: The chatbot makes personal or biographical claims that go beyond light, clearly artificial expressions of personality. The issue applies only when the chatbot refers to having real human experiences, relationships, or memories in a way that blurs its AI identity.
E.6 App Bug Examples
  • Incomplete message: The chatbot’s message is cut off or incomplete, typically due to streaming generation failure, leaving the response unfinished mid-sentence or mid-thought.
  • Transcription errors: Speech-to-text conversion errors cause the chatbot to significantly misunderstand the user, leading to a noticeable impact on the conversation.

### v2

@@
 We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations.
-Starting from Llama-3.3-70B-Instruct as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
+Starting from Llama-3.3-70B-Instruct (grattafiori2024llama3) as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
 $\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(h,c,u)\sim\mathcal{D}}[\log\pi_{\theta}(u|h,c)]$ (1)
@@
 Unlike traditional TOD domains with discrete transactional goals, mental health support offers distinct advantages for studying realistic user simulation: rich emotional dynamics and behavioral complexity, multi-session continuity, diverse naturalistic failure modes beyond binary task success, and production-scale data capturing rare but critical patterns like crisis expressions and resistance to engagement.
-Our dataset consists of anonymized real user conversations with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).
+Our dataset consists of anonymized real user conversations in English with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).
 4.2 Model Configurations
@@
 We perform adversarial training on the final checkpoint (160K), producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations.
-For comparison, we evaluate three zero-shot base models: GPT-4o, Kimi-K2-Instruct, and Llama-3.3-70B-Instruct.
-Training hyperparameters are provided in Appendix C.
+For comparison, we evaluate three zero-shot base models: GPT-4o (openai2024gpt4o), Kimi-K2-Instruct (kimiteam2025k2), and Llama-3.3-70B-Instruct.
+Training hyperparameters are provided in Appendix C.2.
 The discriminators (UserDisc-base, UserDisc-it1, UserDisc-it2, UserDisc-it3) are token classification models trained from the Llama-3.1-8B-Instruct backbone.
@@
 5 Results
+All reported results are from single training runs per configuration.
 5.1 Fine-tuned vs. Zero-shot Comparison
@@
 While our approach demonstrates significant improvements in simulator realism, several limitations remain that warrant future investigation:
+Assumptions and Robustness: Our methodology relies on several key assumptions: (1) discriminator-based rewards accurately reflect true realism rather than exploitable artifacts, (2) the 14 chatbot models used for evaluation are representative of the broader space of possible configurations, and (3) GPT-5-based issue detection accurately identifies problematic conversational features with sufficient reliability to validate simulator quality.
+Results may be sensitive to violations of these assumptions.
 Mode Collapse: While adversarial training substantially improves lexical diversity, fine-tuned simulators still show lower phrase-level variety than real users.
@@
 All data used in this research were collected from real user conversations with our mental health chatbot with explicit user consent.
-All personally identifiable information was removed through anonymization procedures in accordance with data protection policies and ethical guidelines for human subjects research.
+Personally identifiable information was removed through automated script-based detection and removal of names, phone numbers, email addresses, and other identifiers in accordance with data protection policies and ethical guidelines for human subjects research.
 However, due to the sensitive nature of mental health data, we are unable to share the datasets used in this research, despite our desire to promote reproducibility and open science.
+Realistic user simulators could potentially be misused to train systems optimized for engagement over user wellbeing, or to manipulate vulnerable populations.
 Our simulators are designed for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences.
-We note that simulators trained on real conversations may perpetuate biases present in the original dataset; applications should include appropriate oversight and ethical review.
+We note that simulators trained on real conversations may perpetuate biases present in the original dataset, potentially disadvantaging certain demographic groups.
+We emphasize that simulators should be deployed with appropriate oversight and ethical review.
 Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions or edge cases.
+Study procedures were approved as exempt by the Institutional Review Board at New York University School of Medicine (i25-01177).
 Acknowledgments
@@
 This grounds the simulator in the specific topics and emotional tone of the target session, reducing aimless or disengaged responses.
-This session-specific grounding is critical for the simulator to cluster tightly with corresponding real sessions.
 Raw Message History: Message-by-message conversation history from recent sessions, providing detailed behavioral context and immediate conversational flow.
@@
 Figure 4 shows the discrepancy between simulated and real session issue category distributions.
-The simulator over-represents certain categories, including “Chatbot contradicts itself” ($\sim$1.5 percentage points), “Chatbot forgets in context” ($\sim$1.25 percentage points), and “unclear or confusing wording”—likely because simulated users engage more persistently with problematic chatbot behaviors, triggering repeated issues.
-Conversely, the simulator under-represents categories like “duplicate chatbot messages” ($\sim$-2.5 percentage points), “pushing techniques after user request to stop” ($\sim$-2.5 percentage points), and “invalidating or dismissive framing,” which require specific conversational histories that may be less common in simulation.
+The simulator over-represents certain categories, including “Chatbot contradicts itself”, “Chatbot forgets in context”, and “Unclear or confusing wording”—likely because simulated users engage more persistently with problematic chatbot behaviors, triggering repeated issues.
+Conversely, the simulator under-represents categories like “Duplicate chatbot messages”, “Chatbot hallucinates”, and “Premature wrap-up”, which require specific conversational histories that may be less common in simulation.
 The relatively small magnitude of most discrepancies (within $\pm$2.5 percentage points) indicates reasonable distributional alignment overall.
@@
 Critically, no model that failed in simulation improved in real sessions, demonstrating that the simulator successfully identifies ineffective interventions before deployment.
-Appendix C Training Hyperparameters
+Appendix C Model and Training Details
+C.1 Model Licenses and Terms of Use
+All models used in this work are distributed under licenses permitting research use.
+Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct are released under the Llama 3.3 and Llama 3.1 Community License Agreements respectively, which permit commercial and research use with certain restrictions (e.g., services with over 700 million monthly active users require additional licensing).
+GPT-4o is accessed via OpenAI’s API under their Service Terms, which grant users ownership of model outputs for both research and commercial applications.
+Kimi-K2-Instruct is released under a modified MIT license that requires prominent display of the model name when generating content commercially.
+C.2 Training Hyperparameters
 Table 6 shows the hyperparameters used for supervised fine-tuning (SFT) of the base user simulator and Direct Preference Optimization (DPO) during adversarial training iterations.
+C.3 Computational Budget and Infrastructure
+[…]
 Table 6: Hyperparameters for supervised fine-tuning (SFT) and Direct Preference Optimization (DPO).

### v3

@@
-Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization
+DIAL: Direct Iterative Adversarial Learning for Realistic Multi-Turn Dialogue Simulation
 Abstract
-Realistic user simulation is crucial for training and evaluating task-oriented dialogue (TOD) systems, yet creating simulators that accurately replicate human behavior remains challenging.
-A key property of effective simulators is their ability to expose failure modes of the systems they evaluate.
-We present an adversarial training framework that iteratively improves user simulator realism through a competitive dynamic between a generator (user simulator) and a discriminator.
-Applied to mental health support chatbots, our approach demonstrates that fine-tuned simulators dramatically outperform zero-shot base models at surfacing system issues, and adversarial training further enhances diversity, distributional alignment, and predictive validity.
-The resulting simulator achieves a strong correlation between simulated and real failure occurrence rates across diverse chatbot configurations while maintaining low distributional divergence of failure modes.
-Discriminator accuracy decreases drastically after three adversarial iterations, suggesting improved realism.
-These results provide evidence that adversarial training is a promising approach for creating realistic user simulators in mental health support TOD domains, enabling rapid, reliable, and cost-effective system evaluation before deployment.
-Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization
+Realistic user simulation is crucial for training and evaluating multi-turn dialogue systems, yet creating simulators that accurately replicate human behavior remains a significant challenge.
+An effective simulator must expose the failure modes of the systems under evaluation.
+This work introduces Direct Iterative Adversarial Learning (DIAL), a DPO-based adversarial training framework that iteratively enhances user simulator realism through a competitive dynamic between a generator (user simulator) and a discriminator.
+When applied to mental health support, a domain characterized by diverse failure types and a critical dependence on realistic user behavior for failure detection, DIAL restores lexical diversity diminished by supervised fine-tuning and reduces discriminator accuracy from near-perfect to near-random levels.
+The resulting simulator exhibits a strong correlation between simulated and real failure occurrence rates while maintaining low distributional divergence of failure modes.
+These findings indicate that DIAL is a promising method for developing realistic user simulators in multi-turn dialogue, facilitating rapid, reliable, and cost-effective system evaluation prior to deployment.
+DIAL: Direct Iterative Adversarial Learning for Realistic Multi-Turn Dialogue Simulation
 Ziyi Zhu1, Olivier Tieleman1, Caitlin A.
@@
 1 Introduction
-Task-oriented dialogue (TOD) systems guide users through multi-turn conversations to achieve specific goals (schatzmann2006survey).
-The development of large language model (LLM) based TOD systems has created an urgent need for robust evaluation methodologies that go beyond single-turn metrics.
-Multi-turn task-oriented conversations involve complex dynamics where system quality emerges from sustained interaction, user satisfaction accumulates over time, and task success depends on maintaining context across extended dialogue (ouyang2022training; hong2023zeroshotgoaldirecteddialoguerl).
-However, traditional evaluation through live user testing is expensive, time-consuming, and exposes users to potentially suboptimal system behavior.
-User simulation offers a compelling alternative: synthetic agents that replicate human conversational behavior enable rapid iteration, systematic exploration of edge cases, and controlled experimentation (verma2022chai; yu2023promptbased).
-At the current state of the art, realistic user simulators have become essential for training language models through reinforcement learning (RL), where they serve as the environment for policy optimization (snell2023offlinerlnaturallanguage; zhou2024archertraininglanguagemodel).
-Recent advances in multi-turn dialogue optimization have demonstrated the potential of RL methods to improve long-term conversation quality (shani2024multiturnreinforcementlearningpreference; hong2024interactivedialogueagents).
-However, creating simulators that faithfully reproduce the nuanced, diverse, and goal-directed nature of human conversational behavior presents significant challenges.
-Simulators that fail to reproduce realistic user behavior cannot reliably expose system failure modes during development, leading to policies that appear effective in simulation but perform poorly with actual users.
-In this work, we address the challenge of creating realistic user simulators through adversarial training.
-Inspired by Generative Adversarial Networks (GANs) (goodfellow2014generative), we train a discriminator to distinguish between simulated and real user sessions, then use the discriminator feedback to refine the simulator through Direct Preference Optimization (DPO) (rafailov2024direct).
-This creates a competitive dynamic where the simulator learns to fool an increasingly sophisticated discriminator, automatically identifying and correcting the most prominent simulation artifacts.
-We evaluate our approach in the context of Ash [^1] https://www.talktoash.com/, a mental health support chatbot designed by Slingshot AI—a complex TOD domain where sustained conversation quality has direct implications for user wellbeing.
-Our experimental results demonstrate several key findings:
-  • Fine-tuning dramatically outperforms zero-shot: Domain-specific fine-tuning is essential for capturing realistic user behavior distributions in TOD, including negative and resistant patterns that zero-shot models fail to exhibit.
-  • Adversarial training enhances realism: Three iterations of adversarial refinement progressively recover lexical diversity suppressed by supervised learning, increase distributional similarity, and dramatically reduce discriminator accuracy from near-perfect detection to near-random performance.
-  • Strong predictive validity: The best adversarial simulator achieves strong correlation with real model performance and successfully replicates A/B test results on conversation models, enabling reliable offline TOD evaluation before deployment.
-These contributions provide evidence that adversarial training is a promising methodology for creating realistic user simulators in the mental health support TOD domain, with potential applications to RL-based policy optimization and reliable offline evaluation.
-Future work will validate the approach across other TOD domains and integrate simulators into RL training loops.
+Task-oriented dialogue (TOD) systems facilitate multi-turn conversations to help users achieve specific objectives (Schatzmann et al., 2006).
+The emergence of large language models (LLMs) has highlighted the need for robust evaluation methodologies that extend beyond single-turn metrics, as system quality emerges from sustained interaction and task success depends on maintaining context across extended dialogue.
+Traditional evaluation methods, such as live user testing, are costly, time-intensive, and may expose users to suboptimal system behavior.
+User simulation provides a viable alternative, as synthetic agents that mimic human conversational behavior enable rapid iteration, systematic exploration of edge cases, and controlled experimentation (Verma et al., 2022; Yu et al., 2023).
+However, developing simulators that accurately capture the nuanced, diverse, and goal-oriented nature of human conversation remains challenging.
+Simulators that do not reflect realistic user behavior fail to reveal system failure modes during development, resulting in policies that perform well in simulation but poorly with real users.
+This work addresses the challenge of developing realistic user simulators through Direct Iterative Adversarial Learning (DIAL).
+Drawing inspiration from Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), DIAL trains a discriminator to differentiate between simulated and real user sessions, and uses discriminator feedback to refine the simulator via Direct Preference Optimization (DPO) (Rafailov et al., 2023).
+Unlike prior adversarial dialogue approaches that rely on policy gradient methods (Li et al., 2017; Zhao et al., 2021), DIAL avoids training instability and naturally prevents reward hacking through its discrete iteration structure (see Section 3 for details).
+DIAL is evaluated using Ash [^1] https://www.talktoash.com/, a mental health support chatbot developed by Slingshot AI.
+The mental health domain is selected due to its diverse range of failure types, which require realistic user simulation for effective detection of system shortcomings.
+Experimental results reveal several key findings:
+  • DIAL methodology: Three iterations of DIAL refinement using DPO on discriminator-derived preference pairs progressively restore lexical diversity diminished by supervised learning, enhance distributional similarity, and reduce discriminator accuracy from near-perfect to near-random levels.
+  • Failure-sensitive evaluation: Domain-specific fine-tuning, in conjunction with DIAL refinement, is critical for capturing realistic user behaviors, including negative and resistant patterns that zero-shot models do not exhibit.
+This enables accurate identification of diverse chatbot failure modes.
+  • Strong predictive validity: The optimal DIAL-trained simulator demonstrates a strong correlation with real model performance and accurately replicates A/B test results on conversation models, facilitating reliable offline evaluation.
+These findings indicate that DIAL is a promising methodology for developing realistic user simulators in mental health support dialogues, with potential applications in RL-based policy optimization and reliable offline evaluation.
+Future research will extend validation to additional TOD domains and incorporate simulators into end-to-end training processes.
 2 Related Work
 2.1 Multi-Turn Dialogue Optimization
-While RLHF (ouyang2022training) demonstrated that language models can be aligned with human preferences, it typically optimizes single-turn responses rather than multi-turn dialogue outcomes.
-DPO (rafailov2024direct) offers an alternative to RLHF by directly optimizing the policy without explicit reward modeling, simplifying the training pipeline while maintaining effectiveness.
-Optimizing dialogue systems for long-term conversation quality requires methods that go beyond single-turn reward modeling.
-Recent approaches include training on imagined conversations and hindsight regeneration (hong2023zeroshotgoaldirecteddialoguerl; hong2024interactivedialogueagents), token-level policy optimization (shani2024multiturnreinforcementlearningpreference), hierarchical decomposition of planning and execution (zhou2024archertraininglanguagemodel), and offline RL methods for language generation (snell2023offlinerlnaturallanguage).
-These methods critically depend on the quality of the training environment, with realistic user simulators being essential for both policy optimization and evaluation (nam2025efficient).
+Although RLHF (Ouyang et al., 2022) has shown that language models can be aligned with human preferences, its optimization is generally limited to single-turn responses.
+DPO (Rafailov et al., 2023) provides an alternative by directly optimizing the policy without explicit reward modeling, thereby simplifying the training pipeline while preserving effectiveness.
+Enhancing dialogue systems for sustained conversational quality necessitates approaches that extend beyond single-turn reward modeling.
+Recent methods encompass training on imagined conversations and hindsight regeneration (Hong et al., 2023, 2024), token-level policy optimization (Shani et al., 2024), hierarchical decomposition of planning and execution (Zhou et al., 2024), and offline reinforcement learning for language generation (Snell et al., 2023).
+The effectiveness of these approaches is highly contingent on the quality of the training environment, with realistic user simulators being indispensable for both policy optimization and evaluation (Nam et al., 2025).
 2.2 User Simulation for Task-Oriented Dialogue
-User simulation has been central to TOD research since early agenda-based systems that represented behavior through goal-directed dialogue acts (schatzmann2006survey; schatzmann2007agenda).
-Neural approaches later enabled more flexible simulation by capturing dialogue history without rigid structures (elasri2016usersim; kreyssig2018neural), demonstrating that policies trained on neural simulators could achieve higher success rates with real users.
-Recent work on mental health support benchmarking (pombal2024mindeval) demonstrates that LLMs struggle with longer interactions and patients presenting severe symptoms, underscoring the need for realistic simulators that capture the full complexity of human emotions and behaviors.
-LLM-Based Simulation: Large language models have transformed user simulation through in-context learning and prompting approaches (li2022dialogic; davidson2023prompting), achieving goal success rates comparable to human-bot interactions with greater linguistic diversity and minimal seed dialogues.
-However, prompting-only approaches suffer from hallucinations and goal drift.
-Fine-tuning addresses these limitations through various techniques, including reducing behavioral inconsistencies, joint policy and NLG optimization for domain transfer, and dual-model architectures with chain-of-thought reasoning (sekulic2024daus; lin2022gentus; luo2024duetsim).
-RL-Based Simulator Training: While supervised learning dominates simulator training, reinforcement learning approaches optimize for task-level objectives rather than imitation alone.
-Multi-agent RL enables joint system-user training (takanobu2020guided; liu2018dialogue), while adversarial training creates competitive dynamics between simulators and discriminators (keizer2023adversarial; li2017adversarial).
-The adversarial approach has demonstrated substantial improvements, with discriminator-based rewards being "more general, natural and effective" than hand-crafted metrics.
-Our work extends this adversarial paradigm to LLM-based simulators, using DPO rather than policy gradient methods to leverage discriminator feedback while avoiding reward hacking.
-Evaluation Methodology: Traditional TOD evaluation relies on task-level metrics such as success rate and slot error rate measured on standard benchmarks like MultiWOZ (budzianowski2018multiwoz), SGD (rastogi2020schema), and tau-bench (yao2024taubench), which effectively capture transactional outcomes but provide limited insight into conversational quality.
-The poor correlation between surface-level metrics like BLEU/ROUGE and human judgments (liu2016not) motivates more sophisticated evaluation approaches that consider behavioral fidelity and distributional alignment (nam2025efficient).
-Our work advances this evaluation paradigm through three key contributions: (1) explicit modeling of conversational feature occurrence rate distributions using KL divergence to measure alignment on fine-grained failure modes rather than binary task success, (2) validation of predictive utility through model A/B test replication beyond correlation analysis, and (3) discriminator accuracy as a direct measure of distributional realism.
+User simulation has played a central role in TOD research since the development of early agenda-based systems, which modeled behavior through goal-directed dialogue acts (Schatzmann et al., 2006, 2009).
+Subsequent neural approaches enabled more flexible simulation by capturing dialogue history without rigid structural constraints (El Asri et al., 2016; Kreyssig et al., 2018), and demonstrated that policies trained on neural simulators can achieve higher success rates with real users.
+Recent benchmarking in mental health support (Pombal et al., 2025) reveals that LLMs encounter difficulties with extended interactions and patients exhibiting severe symptoms, highlighting the necessity for realistic simulators that reflect the full complexity of human emotions and behaviors.
+LLM-Based Simulation: LLMs have significantly advanced user simulation by leveraging in-context learning and prompting methods (Li et al., 2022; Davidson et al., 2024), achieving goal success rates comparable to human-bot interactions while offering greater linguistic diversity and requiring minimal seed dialogues.
+Nevertheless, prompting-only strategies are prone to hallucinations and goal drift.
+Fine-tuning mitigates these issues through techniques such as reducing behavioral inconsistencies, jointly optimizing policy and natural language generation (NLG) for domain transfer, and employing dual-model architectures with chain-of-thought reasoning (Sekulić et al., 2024; Lin et al., 2022; Luo et al., 2024).
+RL-Based Simulator Training: While supervised learning dominates simulator training, reinforcement learning methods optimize for task-level objectives beyond mere imitation.
+Multi-agent reinforcement learning facilitates joint system-user training (Takanobu et al., 2020; Liu and Lane, 2017), while adversarial training introduces competitive dynamics between simulators and discriminators (Keizer et al., 2023; Li et al., 2017; Zhao et al., 2021).
+Li et al.
+(2017) employ REINFORCE with Monte Carlo search to backpropagate discriminator rewards, but report that this leads to unstable training and necessitates teacher forcing as a regularizer.
+Zhao et al.
+(2021) also utilize GAN-based policy gradient methods for user simulation in recommendation systems.
+The present work contributes methodologically by extending the adversarial paradigm to LLM-based simulators through DIAL, with a key distinction in optimization: DPO is applied to discriminator-derived preference pairs instead of policy gradient methods.
+DIAL circumvents the instability associated with policy gradients in text generation and inherently prevents reward hacking (see Section 3).
+Evaluation Methodology: Traditional TOD evaluation, including recent end-to-end frameworks such as AutoTOD (Xu et al., 2024) and ProTOD (Dong et al., 2025), relies on task-level metrics such as success rate and slot accuracy measured on standard benchmarks like MultiWOZ (Budzianowski et al., 2018), SGD (Rastogi et al., 2020), and tau-bench (Yao et al., 2024).
+These metrics effectively capture transactional outcomes but provide limited insight into conversational quality.
+Our work addresses the complementary problem of building realistic user simulators to evaluate such systems, focusing on failure mode detection rather than transactional success, and advances this evaluation paradigm through three key contributions: (1) explicit modeling of conversational feature occurrence rate distributions using KL divergence to measure alignment on fine-grained failure modes, (2) validation of predictive utility through model A/B test replication beyond correlation analysis, and (3) discriminator accuracy as a direct measure of distributional realism.
 3 Methodology
@@
 A critical design choice for realistic user simulation is the construction of context $c$ that conditions the simulator’s behavior.
-Effective context should provide: (1) role and behavioral guidelines through system instructions, (2) user background and goals, (3) conversational history to maintain consistency, and (4) session-specific grounding to anchor the simulator to target conversation characteristics.
+Effective context should provide: (1) role and behavioral guidelines, (2) user background and goals, (3) conversational history to maintain consistency, and (4) task-specific grounding to anchor the simulation.
 The specific context structure depends on the domain—transactional TOD systems might include task schemas and slot values, while open-ended domains might emphasize user profiles and session summaries.
-Our mental health support chatbot implementation (detailed in Appendix A) uses a hierarchical structure spanning user profiles, multi-session summaries, and current session grounding.
-As we demonstrate in Section 5, rich context grounding enables fine-tuned simulators to cluster tightly with their corresponding real sessions in embedding space, while zero-shot models without such grounding form distinct clusters, indicating systematic behavioral drift.
+Our mental health support chatbot (detailed in Appendix A) uses a hierarchical structure spanning user profiles, multi-session summaries, and current session grounding.
+As we demonstrate in Section 5, rich context grounding enables fine-tuned simulators to cluster tightly with their corresponding real sessions in embedding space, while zero-shot models exhibit systematic behavioral drift.
 3.3 Base Simulator Training
 We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations.
-Starting from Llama-3.3-70B-Instruct (grattafiori2024llama3) as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
+Starting from Llama-3.3-70B-Instruct (Grattafiori et al., 2024) as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
 $\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(h,c,u)\sim\mathcal{D}}[\log\pi_{\theta}(u|h,c)]$ (1)
@@
 To measure simulator realism, we train a discriminator $D_{\phi}$ to distinguish between simulated and real user sessions.
-We implement the discriminator as a token classification model built on the Llama-3.1-8B-Instruct backbone.
-The discriminator processes the full conversation and predicts after each user message whether the session is real or simulated:
+We implement the discriminator as a token classification model which processes the full conversation and predicts after each user message whether the session is real or simulated:
 $p_{t}=D_{\phi}(\text{real}|h_{t},c)$ (2)
 We use causal attention in the discriminator, meaning predictions are made based on the current message and all previous messages in the session.
-This design choice is essential because any individual message could plausibly be either real or simulated when taken out of context.
-The discriminator’s effectiveness stems from detecting distributional patterns across the conversation: certain phrases repeated multiple times, unnatural topic progressions, or behavioral patterns that diverge from real user dynamics.
+The discriminator’s effectiveness stems from detecting distributional patterns across the conversation: repeated phrases, unnatural topic progressions, or behavioral patterns that diverge from real user dynamics.
 Without conversational context, the discrimination task becomes noisy and ill-defined, as isolated messages lack the sequential dependencies that reveal simulation artifacts.
@@
 A high discriminator accuracy indicates the simulator has not achieved realistic behavior; our goal is to train simulators that reduce discriminator accuracy toward random guessing.
-In our adversarial training approach, each new discriminator is trained exclusively on data from the most recent simulator version.
+In DIAL, each new discriminator is trained exclusively on data from the most recent simulator version.
 This ensures the preference pairs we generate for DPO reflect the current policy’s behavior rather than artifacts from earlier iterations.
-Research on preference optimization supports this on-policy approach: on-policy preference pairs exhibit stronger contrastiveness and clearer preference signals (zhao2024reflective), online methods require only “partial coverage” compared to the strict “global coverage” needed by offline methods (song2024importance), and on-policy sampling efficiently concentrates probability mass on high-reward responses (tajwar2024preference).
-3.5 Adversarial Training Framework
-Our adversarial training process follows an iterative refinement loop, illustrated in Figure 1. Algorithm 1 provides the detailed procedure.
-Algorithm 1 Adversarial User Simulator Training
+Research on preference optimization supports this on-policy approach: on-policy preference pairs exhibit stronger contrastiveness and clearer preference signals (Zhao and Li, 2025), online methods require only “partial coverage” compared to the strict “global coverage” needed by offline methods (Song et al., 2024), and on-policy sampling efficiently concentrates probability mass on high-reward responses (Tajwar et al., 2024).
+3.5 DIAL: Direct Iterative Adversarial Learning
+The DIAL training process follows an iterative refinement loop, illustrated in Figure 1. Algorithm 1 provides the detailed procedure.
+Algorithm 1 DIAL: Direct Iterative Adversarial Learning
 1: Initialize user simulator $\pi_{\theta}$ from SFT
@@
 10:end for
-Figure 1: Overview of the adversarial training process.
+Figure 1: Overview of the DIAL training process.
 The simulator generates sessions that are evaluated by the discriminator to distinguish real from simulated conversations.
 Discriminator feedback guides preference dataset generation, which is then used to refine the simulator through DPO, creating an iterative competitive dynamic that improves realism.
-Optimization Method: We use DPO rather than RLHF methods (schulman2017proximal; shao2024deepseekmath) because discriminator rewards are inherently policy-dependent and vulnerable to reward hacking.
+Optimization Method: DIAL uses DPO rather than RLHF methods (Schulman et al., 2017; Shao et al., 2024) because discriminator rewards are inherently policy-dependent and vulnerable to reward hacking.
 With RLHF, the simulator would continuously optimize against a fixed reward model, eventually exploiting patterns that score highly but don’t reflect genuine realism.
@@
 Higher rewards indicate messages that make the session appear more realistic.
-Preference Dataset Generation: At messages with the highest and lowest rewards (critical moments where the discriminator most clearly detects unrealism), we sample 8 alternative responses.
+Preference Dataset Generation: At messages with the highest and lowest rewards, we sample 8 alternative responses.
 These alternatives are ranked by their discriminator-based rewards to create preference pairs $(u_{\text{chosen}},u_{\text{rejected}})$ where the chosen response has the highest reward.
@@
 This selects high-confidence pairs with maximum learning signal while preventing over-correction.
-DPO Update: We train the new simulator using Direct Preference Optimization (rafailov2024direct), which directly optimizes the policy to prefer higher-reward responses.
+DPO Update: We train the new simulator using Direct Preference Optimization (Rafailov et al., 2023), which directly optimizes the policy to prefer higher-reward responses.
 We use the previous iteration’s simulator as the reference policy $\pi_{\text{ref}}$.
@@
 4.1 Domain and Data
-We evaluate our methodology in the context of mental health support—a complex task-oriented dialogue domain where users pursue mental health goals (emotional regulation, insight development, behavioral change) through extended multi-turn conversations.
-Unlike traditional TOD domains with discrete transactional goals, mental health support offers distinct advantages for studying realistic user simulation: rich emotional dynamics and behavioral complexity, multi-session continuity, diverse naturalistic failure modes beyond binary task success, and production-scale data capturing rare but critical patterns like crisis expressions and resistance to engagement.
-Our dataset consists of anonymized real user conversations in English with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).
+DIAL is evaluated within the domain of mental health support, in which users pursue mental health goals such as emotional regulation, insight development, and behavioral change through extended multi-turn conversations.
+This domain was selected to specifically assess simulators’ capacity to reveal failure modes of TOD systems, as mental health conversations present a particularly diverse array of failure types (see Appendix E).
+In contrast to traditional TOD benchmarks such as MultiWOZ (Budzianowski et al., 2018) and SGD (Rastogi et al., 2020), where failure modes are primarily transactional (e.g., incorrect slot filling or task incompletion) and less reliant on realistic user simulation, the mental health support domain features complex emotional dynamics and naturalistic failure modes that require faithful user behavior for effective detection.
+Furthermore, production-scale data in this domain captures rare but critical phenomena, such as crisis expressions and resistance to engagement, thereby providing a rigorous testbed for evaluating simulator realism.
+Our dataset consists of anonymized real user conversations in English with our mental health support chatbot, including session-level context.
 4.2 Model Configurations
-We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (hu2022lora) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects.
-We perform adversarial training on the final checkpoint (160K), producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations.
-For comparison, we evaluate three zero-shot base models: GPT-4o (openai2024gpt4o), Kimi-K2-Instruct (kimiteam2025k2), and Llama-3.3-70B-Instruct.
+We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (Hu et al., 2022) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects.
+We apply DIAL on the final checkpoint, producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations.
+For comparison, we evaluate three zero-shot base models: GPT-4o (Hurst et al., 2024), Kimi-K2-Instruct (Bai et al., 2025), and Llama-3.3-70B-Instruct.
 Training hyperparameters are provided in Appendix C.2.
@@
 For each configuration, we generate 2,000 simulated sessions: 1,000 unique session contexts drawn from real user profiles and conversation histories × 2 simulations per context, with early stopping when conversations naturally conclude.
-Our context construction for user simulation follows the hierarchical structure described in Appendix A, including user profiles, multi-session history, and current session grounding.
 To isolate simulator quality effects, we use a fixed set of 14 chatbot models to simulate conversations.
@@
 Higher entropy indicates more diverse word usage.
-Type-Token Ratio (TTR): The ratio of unique words (types) to total words (tokens) in a session, measuring lexical diversity.
-TTR ranges from 0 to 1, with higher values indicating greater vocabulary variety.
+Type-Token Ratio (TTR): The ratio of unique words (types) to total words (tokens) in a session, measuring lexical diversity. Higher TTR indicates greater vocabulary variety.
 Bigram Diversity: The ratio of unique word pairs (bigrams) to total bigrams in a session, measuring phrase-level variety and reducing sensitivity to repetitive phrasing.
 4.3.2 Behavioral Metrics
-User Action Distribution: Conversations are annotated with mental health-relevant user actions using GPT-5-nano to classify each individual user message according to our action taxonomy (see Appendix D for example categories).
-Since individual messages can have multiple actions, the action distributions are not normalized (probabilities do not sum to 1).
+User Action Distribution: Conversations are annotated with task-relevant user actions using GPT-5-nano to classify each individual user message according to our action taxonomy (see Appendix D).
+Since individual messages can have multiple actions, the action distributions are not normalized.
 Issue Rate Analysis: A critical requirement for realistic user simulation is the ability to expose chatbot failure modes.
 We evaluate simulators by measuring how well they replicate the rate and distribution of problematic conversational features (“issues”) that chatbots exhibit with real users.
-Issues are identified using GPT-5 with low reasoning effort to analyze entire sessions.
-The analyzer identifies issues from a detailed taxonomy (35+ categories including invalidating responses, unclear wording, hallucinations, etc.) and highlights the specific messages where each issue occurred, enabling proper credit assignment (see Appendix E for example categories).
-We compare issue rates and category distributions between simulated and real conversations to validate that our simulator accurately reflects the chatbot’s real-world performance and can reliably identify system weaknesses before deployment.
-Since we examine proportions of issue categories (which form a valid probability distribution that sums to 1), we use KL divergence to measure distributional alignment:
+Issues are identified using GPT-5 with low reasoning effort.
+The analyzer identifies issues from a detailed taxonomy (35+ categories including invalidating responses, unclear wording, hallucinations, etc.) and highlights the specific messages where each issue occurred, enabling proper credit assignment (see Appendix E).
+We compare issue rates and distribution between simulated and real conversations to validate that our simulator accurately reflects the chatbot’s real-world performance and can reliably identify system weaknesses before deployment.
+We use KL divergence to measure distributional alignment:
 $\text{KL}(p\|q)=\sum_{i}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)$ (6)
 where $p$ is the real distribution and $q$ is the simulated distribution.
-Similar to the discriminator, the issue analyzer examines whole sessions rather than isolated messages.
-This design is essential because messages are only problematic when considered in context with the user’s situation and their subsequent replies.
-The session-level approach makes realistic user simulation critical for accurate evaluation—the simulator must produce contextually appropriate responses that reveal when the chatbot’s messages fail to address the user’s actual needs.
-This enables numerous evaluation dimensions that would be impossible with isolated message analysis: detecting circular conversations, identifying premature topic changes, recognizing when the chatbot ignores user requests, and assessing whether mental health progress occurs naturally over the conversation arc.
+Similar to the discriminator, the issue analyzer examines whole sessions rather than isolated messages, which are only problematic when considered in context with the user’s situation and their subsequent replies.
+This enables numerous evaluation dimensions that would be impossible with isolated message analysis: detecting circular conversations, identifying premature topic changes, and assessing whether mental health progress occurs naturally over the conversation arc.
 4.3.3 Discriminator Performance
-We measure discriminator accuracy using fine-tuned discriminators trained on balanced datasets of real and simulated sessions from their corresponding simulator versions.
-Each discriminator is trained under identical conditions—1,600 training sessions and 400 validation sessions, with the same learning rate and number of epochs—ensuring that accuracy differences reflect simulator quality rather than training differences.
-Lower discriminator accuracy indicates a more realistic simulator, with the target being 50% (random guessing).
+We measure discriminator accuracy using fine-tuned discriminators trained on balanced datasets of real and simulated sessions from their corresponding simulators.
+Each discriminator is trained under identical conditions—1,600 training sessions and 400 validation sessions, with the same hyperparameters—ensuring that accuracy differences reflect simulator quality rather than training differences.
+Lower discriminator accuracy indicates a more realistic simulator.
 5 Results
@@
 5.1 Fine-tuned vs. Zero-shot Comparison
-Fine-tuned simulators dramatically outperform zero-shot models across all metrics.
+As a foundation for DIAL, we first establish that domain-specific SFT is necessary for realistic simulation.
+Fine-tuned simulators substantially outperform zero-shot models across all metrics.
 t-SNE visualization (Appendix B.1) shows that embeddings of conversations simulated by base models form distinct clusters separate from their real counterparts.
-The fine-tuned simulators’ outputs, on the other hand, form tight clusters defined by shared context rather than by the generation process, demonstrating effective grounding.
-Zero-shot models produce unrealistically compliant users, as evidenced by significantly lower issue rates (Figure 2) and severe overrepresentation of positive mental health indicators (Appendix B.3).
-Table 1: Linguistic diversity features comparing real users with simulators.
-Entropy measures vocabulary diversity; TTR (type-token ratio) measures lexical diversity; Diversity measures bigram diversity.
-Values closer to real users indicate better linguistic realism.
-Best fine-tuned values (closest to real users) are in bold.
+The fine-tuned simulators’ outputs, on the other hand, form tight clusters defined by shared context, demonstrating effective grounding.
+Zero-shot models produce unrealistically compliant users, as evidenced by significantly lower issue rates (Figure 2) and overrepresentation of positive mental health indicators (Appendix B.3).
+Table 1: Linguistic diversity features comparing real users with simulators. Values closer to real users indicate better linguistic realism. Best fine-tuned values (closest to real users) are in bold.
 Simulator Entropy TTR Bigram
@@
 UserSim-160K-it3 6.51 0.461 0.841
-As shown in Table 1, which compares linguistic features across simulators, some base models achieve diversity metrics closer to real users than fine-tuned models (e.g., Kimi-K2-Instruct’s bigram diversity of 0.851 vs.
-real users’ 0.861) as supervised fine-tuning exhibits non-monotonic effects on lexical diversity.
-Bigram and TTR initially improve from 10K-40K samples as the model learns to use conversational context to improve response diversity.
+As shown in Table 1, which compares linguistic features across simulators, some base models achieve diversity metrics closer to real users than fine-tuned models as SFT exhibits non-monotonic effects on lexical diversity.
+Bigram and TTR initially improve from 10K-40K samples as the model learns to use conversational context to improve diversity.
 However, mode collapse emerges from 40K-160K samples as continued training learns common patterns at the expense of rare variations.
-For behavioral alignment, fine-tuned simulators dramatically outperform zero-shot base models, achieving total BCE losses of .159-.160 compared to .187-.207 for zero-shot models (see Appendix B.3).
-To evaluate how well simulators surface chatbot issues, we measure issue rates (percentage of chatbot messages containing issues) using GPT-5 to analyze complete sessions.
+For behavioral alignment, fine-tuned simulators dramatically outperform zero-shot base models (see Appendix B.3).
+To evaluate how well simulators surface chatbot issues, we measure issue rates as the percentage of chatbot messages containing issues.
 Figure 2 shows that fine-tuned simulators produce rates of 3.07-3.31%, closely matching real conversations (3.02%), while zero-shot models show substantially lower rates of 0.98-2.54%.
-UserSim-160K-it2 achieves the closest calibration to the real rate (3.07%).
 Detailed issue category discrepancies are provided in Appendix B.2.
@@
 Zero-shot base models significantly underestimate issue rates, while fine-tuned models achieve realistic calibration.
-5.2 Adversarial Training Improvements
-Three iterations of adversarial training produced progressive improvements:
+5.2 DIAL Training Improvements
+Three iterations of DIAL produced progressive improvements:
 Table 2: Discriminator performance (accuracy, F1, MCC) at different conversation turns.
@@
 UserDisc-it3 52.5 60.6 5.5 64.0 64.4 28.0 68.1 59.5 33.9 84.5 78.9 67.3 69.5 63.9 37.6
-Linguistic Diversity: Adversarial training successfully addresses mode collapse by recovering lexical diversity while maintaining behavioral alignment.
-UserSim-160K-it3 progressively approaches real user levels across all diversity metrics, demonstrating that adversarial training addresses a fundamental limitation of supervised fine-tuning.
-However, oscillatory behavior across iterations indicates the need for careful hyperparameter tuning.
-We employ conservative dataset filtering (retaining only high-confidence preference pairs) to mitigate over-correction and maintain distribution stability.
-Distributional Alignment: Adversarial training produces modest improvements in action distribution alignment (see Appendix B.3), with progressive refinement across iterations.
-The limited magnitude of improvements suggests the base fine-tuned model already achieved strong behavioral alignment.
-Discriminator Fooling: Adversarial training dramatically reduced discriminator accuracy across conversation turns (Table 2).
-All discriminators were trained under identical conditions with on-policy optimization—each trained exclusively on sessions from its corresponding simulator version.
-The dramatic performance degradation confirms that adversarial iterations remove systematic artifacts and produce strong distributional alignment throughout full sessions.
+Linguistic Diversity: DIAL successfully addresses mode collapse by recovering lexical diversity while maintaining behavioral alignment, as models progressively approach real user levels across all diversity metrics.
+However, non-monotonic improvement across iterations indicates the need for careful hyperparameter tuning.
+We employ conservative dataset filtering to mitigate over-correction and maintain stability.
+Distributional Alignment: DIAL produces modest improvements in action distribution alignment (see Appendix B.3).
+The limited magnitude of improvements suggests the base fine-tuned model already achieved strong alignment.
+Discriminator Fooling: DIAL dramatically reduced discriminator accuracy across conversation turns (Table 2).
+All discriminators were trained under identical conditions with on-policy optimization.
+The dramatic performance degradation confirms that DIAL iterations remove systematic artifacts and produce strong distributional alignment throughout sessions.
 Predictive Validity: To evaluate how well simulators enable reliable offline evaluation, we measure correlation between simulated and real issue rates across diverse chatbot models, and KL divergence between simulated and real issue category distributions.
-Table 3 shows that adversarial training produces substantial improvements in both metrics.
+Table 3 shows that DIAL produces substantial improvements in both metrics.
 UserSim-160K-it2 achieves the strongest correlation ($r=0.818$), while UserSim-160K-it3 achieves the lowest distributional divergence (KL = 0.148).
-This provides compelling evidence that adversarially trained simulators are significantly more realistic and effective at identifying chatbot problems in offline testing compared to their base fine-tuned counterparts.
-Notably, aggregating predictions across all simulator models achieves a strong correlation ($r=0.801$) and low distributional divergence (KL = 0.152), suggesting that ensembling multiple simulators could provide even more robust evaluation.
+This provides compelling evidence that DIAL-trained simulators are significantly more realistic and effective at identifying chatbot problems in offline testing compared to their base fine-tuned counterparts.
+Notably, aggregating predictions across all simulator models achieves both a strong correlation and low distributional divergence, suggesting that ensembling multiple simulators could provide even more robust evaluation.
 Table 3: Correlation with real issue rates and KL divergence of issue category distributions.
@@
 Aggregate 0.801 0.0006 0.152
-A/B Test Replication: Beyond aggregate correlation, we validate predictive utility by replicating a recent model A/B test.
+A/B Test Replication: Beyond correlation, we validate predictive utility by replicating a recent model A/B test.
 All three models exhibited consistent directional trends between simulation and real data, and no model that failed in simulation improved in the wild (Appendix B.5).
-Combined with strong correlation and distributional alignment, these results validate the simulator’s utility for reliable offline evaluation.
 6 Discussion
 Our results establish several important findings for user simulator development:
-Firstly, zero-shot models exhibit severe distributional biases, producing unrealistically compliant users who systematically underrepresent negative and resistant behaviors.
-This confirms that generic language models, despite their capabilities, cannot reliably replicate domain-specific user behavior without exposure to real conversational patterns.
-While supervised fine-tuning achieves strong behavioral alignment, it suppresses lexical diversity through mode collapse.
-Adversarial refinement recovers this diversity while maintaining behavioral fidelity—successfully producing simulators that fool increasingly sophisticated discriminators without sacrificing predictive utility.
-This demonstrates that realism requires both distributional accuracy and surface-level variation.
-Lastly, unlike traditional metrics that measure task success or correlation, discriminator accuracy directly quantifies distributional realism by testing whether conversations are distinguishable from real interactions.
-The dramatic accuracy degradation across adversarial iterations provides evidence of improved simulation fidelity that complements task-level validation.
+As context for the substantive contributions, note that zero-shot models show severe distributional biases, producing unrealistically compliant users who underrepresent negative and resistant behaviors.
+This finding confirms that domain-specific fine-tuning is a necessary foundation for realistic simulation, consistent with the broader literature on domain adaptation.
+While SFT achieves strong behavioral alignment, it suppresses lexical diversity through mode collapse.
+DIAL, which avoids the training instability of prior policy gradient approaches (Li et al., 2017; Zhao et al., 2021) by employing a discrete iteration structure and preference-based optimization, recovers this diversity while maintaining behavioral fidelity.
+DIAL produces simulators that can deceive increasingly sophisticated discriminators, demonstrating that realism requires both distributional accuracy and surface-level variation.
+Unlike traditional metrics that measure task success or correlation, discriminator accuracy directly quantifies distributional realism by testing whether conversations are distinguishable from real interactions.
 7 Conclusion
-We have presented an adversarial training framework for creating realistic user simulators that can reliably expose failure modes in multi-turn task-oriented dialogue systems.
-Applied to mental health support—a complex TOD domain where users pursue mental health goals through extended goal-directed conversations—our approach demonstrates that domain-specific fine-tuning combined with adversarial refinement produces simulators that closely replicate real user behavior across linguistic, semantic, and behavioral dimensions.
-Our key contributions are: (1) demonstration that fine-tuned models vastly outperform zero-shot base models for behavioral fidelity, (2) an adversarial training method that recovers lexical diversity suppressed by supervised learning while improving distributional alignment and predictive validity, and (3) comprehensive validation that adversarially-trained simulators reliably predict real model performance, achieving strong correlation with real-world results and successfully replicating A/B test trends.
-These results provide evidence that adversarial training is a promising methodology for creating realistic user simulators in mental health support TOD domains.
-By enabling rapid iteration and systematic exploration of system failure modes before deployment, such simulators can reduce development time, protect users from suboptimal experiences, and potentially provide safe environments for reinforcement learning based optimization of conversation quality.
+This work presents DIAL (Direct Iterative Adversarial Learning), a framework for creating realistic user simulators that can reliably expose failure modes in multi-turn dialogue systems.
+Applied to mental health support, DIAL shows that domain-specific fine-tuning combined with iterative adversarial refinement produces simulators that closely replicate real user behavior across linguistic, semantic, and behavioral dimensions.
+The key contributions of our work are: (1) the DIAL methodology, which uses discriminator-derived preference pairs with DPO instead of policy gradient methods, avoiding the training instability of prior adversarial dialogue approaches while recovering lexical diversity suppressed by supervised learning and improving distributional alignment; (2) demonstration that DIAL enables accurate identification of diverse failure modes in the mental health support domain; and (3) comprehensive validation showing that DIAL-trained simulators reliably predict real model performance, achieving strong correlation with real-world results and replicating A/B test trends.
+These results show that DIAL is a promising methodology for creating realistic user simulators in multi-turn dialogue systems.
+By enabling rapid iteration and systematic exploration of system failure modes before deployment, such simulators can reduce development time, protect users from suboptimal experiences, and offer safe environments for RL-based optimization of conversation quality.
 Limitations
@@
 Results may be sensitive to violations of these assumptions.
-Mode Collapse: While adversarial training substantially improves lexical diversity, fine-tuned simulators still show lower phrase-level variety than real users.
-Techniques like controlled generation or mixture-of-experts architectures might further improve diversity preservation.
-Domain Transfer: Our evaluation focuses on mental health support—a complex, open-ended TOD domain.
-Validating the methodology in traditional TOD domains (restaurant booking, customer service) and other complex domains (tutoring, technical support) would establish broader applicability and identify domain-specific challenges.
-RL Training Integration: While our simulators enable valid offline evaluation, integrating them directly into RL training loops for TOD policy optimization remains future work.
-Recent work on RL-based tutoring (nam2025efficient) demonstrates that simulator quality is critical for reliable RL-based dialogue optimization, underscoring the importance of adversarial training for creating simulators that support both evaluation and training.
+Mode Collapse: Although DIAL substantially improves lexical diversity, fine-tuned simulators still exhibit lower phrase-level variety than real users.
+Techniques such as controlled generation or mixture-of-experts architectures may further enhance diversity preservation.
+Domain Transfer: The evaluation in this study focuses on mental health support, a complex and open-ended TOD domain.
+Validating the methodology on traditional TOD benchmarks such as MultiWOZ (Budzianowski et al., 2018) and SGD (Rastogi et al., 2020), as well as in other complex domains such as tutoring and technical support, would establish broader applicability and help identify domain-specific challenges.
+End-to-End TOD Training and Evaluation: Although the current simulators enable valid offline evaluation of chatbot failure modes, their application for training and evaluating end-to-end dialogue policies remains an area for future investigation.
+It is hypothesized that increasing simulator realism may not necessarily improve raw task success rates when evaluated by the simulator itself.
+Instead, success rates are expected to become more calibrated with real-user outcomes, as less realistic simulators often exhibit excessive compliance, leading to inflated success metrics.
+For instance, Cheng et al.
+(2022) train a user simulator with RL using task success as the reward, which encourages the simulator to become more compliant, contrary to the goal of surfacing system failures.
+Rigorous validation of this hypothesis requires access to real users for ground-truth comparison, a resource that is challenging to obtain at scale for most TOD benchmarks.
+Human Evaluation: The evaluation of simulator realism in this study relies on automated metrics (discriminator accuracy, distributional alignment, predictive validity) rather than direct human judgment.
+Complementing these metrics with human evaluation of simulated conversation realism would provide additional validation and help identify failure modes that automated metrics may overlook.
 Ethical Considerations
-All data used in this research were collected from real user conversations with our mental health chatbot with explicit user consent.
+All data used in this research were collected from real user conversations with the mental health chatbot, with explicit user consent.
 Personally identifiable information was removed through automated script-based detection and removal of names, phone numbers, email addresses, and other identifiers in accordance with data protection policies and ethical guidelines for human subjects research.
-However, due to the sensitive nature of mental health data, we are unable to share the datasets used in this research, despite our desire to promote reproducibility and open science.
-Realistic user simulators could potentially be misused to train systems optimized for engagement over user wellbeing, or to manipulate vulnerable populations.
-Our simulators are designed for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences.
-We note that simulators trained on real conversations may perpetuate biases present in the original dataset, potentially disadvantaging certain demographic groups.
-We emphasize that simulators should be deployed with appropriate oversight and ethical review.
-Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions or edge cases.
+However, due to the sensitive nature of mental health data, the datasets used in this research cannot be shared, despite the intention to promote reproducibility and open science.
+Realistic user simulators could potentially be misused to train systems optimized for engagement rather than user wellbeing, or to manipulate vulnerable populations.
+The simulators developed in this work are intended for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences.
+It should be noted that simulators trained on real conversations may perpetuate biases present in the original dataset, potentially disadvantaging certain demographic groups.
+Simulators should therefore be deployed with appropriate oversight and ethical review.
+Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions.
 Study procedures were approved as exempt by the Institutional Review Board at New York University School of Medicine (i25-01177).
@@
 $\sim$15% in real data) and “Agrees with Chatbot.” This creates an artificially positive mental health support environment that fails to challenge the chatbot or surface issues.
-Fine-tuned simulators (UserSim variants) achieve close alignment with real sessions across both indicator types, with adversarial training iterations showing incremental improvements in matching real distributions.
+Fine-tuned simulators (UserSim variants) achieve close alignment with real sessions across both indicator types, with DIAL iterations showing incremental improvements in matching real distributions.
 We use Binary Cross-Entropy (BCE) to measure distributional alignment:
@@
 It is calculated using a t-test: $t=r\sqrt{(n-2)/(1-r^{2})}$ with n2n-2 degrees of freedom.
-The adversarial training process substantially improves correlation with real issue rates, as shown in Table 3.
-UserSim-160K-it2 achieves $r=0.818$ compared to UserSim-160K-base’s $r=0.528$, demonstrating the value of adversarial refinement.
-However, we note that several models achieve strong correlations even without adversarial training.
+DIAL substantially improves correlation with real issue rates, as shown in Table 3.
+UserSim-160K-it2 achieves $r=0.818$ compared to UserSim-160K-base’s $r=0.528$, demonstrating the value of DIAL refinement.
+However, we note that several models achieve strong correlations even without DIAL.
 This suggests that for rolled-out evaluation or RL training applications, ensembling multiple user simulator models could reduce noise and improve stability.
@@
 C.2 Training Hyperparameters
-Table 6 shows the hyperparameters used for supervised fine-tuning (SFT) of the base user simulator and Direct Preference Optimization (DPO) during adversarial training iterations.
+Table 6 shows the hyperparameters used for SFT of the base user simulator and DPO during DIAL iterations.
 C.3 Computational Budget and Infrastructure
@@
 This appendix provides example user action categories from our taxonomy used to annotate and analyze user behavior in mental health support conversations.
-Each user message is classified using GPT-5-nano.
 Actions are classified as positive mental health indicators (Good), negative indicators (Bad), or user features that should change over time.
@@
 This appendix provides example issue categories from our taxonomy used to evaluate chatbot quality.
-Issues are identified by GPT-5 analyzing entire conversation sessions with low reasoning effort.
-For each detected issue, the analyzer highlights the specific affected messages to enable proper credit assignment.
 The full taxonomy contains 35+ categories organized into six main categories (User Distress, User Confusion, User Frustration, User Losing Trust, Chatbot Violates Policies, App Bugs); we highlight key examples below.